A polyhedron with 6 vertices and 12 edges is shown below. What is the total number of faces for this polyhedron?
Answer by AnlytcPhil(1806)

We use Euler's polyhedron formula 

V -  E + F = 2
6 - 12 + F = 2
    -6 + F = 2
         F = 8

It has 8 faces.

Edwin